OpenStudy (anonymous):

Can sombody help me find the solution? 1/2x^2 – x + 5 = 0

OpenStudy (anonymous):

@zzr0ck3r ??

OpenStudy (zzr0ck3r):

1/(2x^2) - x + 5 = 0 ?

OpenStudy (anonymous):

the equation is 1halfx^2-x+5=0

OpenStudy (zzr0ck3r):

(1/2)x^2 - x + 5 = 0?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

= (1/2)(x^2 - 2x + 10) = 0 (1/2)(x-1)^2 = -5+1/2 (x-1)^2 = -9 x-1 = +-sqrt(-9) = +-3i x = -3i+1 x = 3i+1

OpenStudy (zzr0ck3r):

I factored the 1/2 out then completed the square

OpenStudy (zzr0ck3r):

you can use this also (-b+-sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

These are the multiple choice answers: a) 1 ± i sqrt 11 b) –1 ± i sqrt 9 c) 1 ± i sqrt 9 d) –1 ± i sqrt 11

OpenStudy (anonymous):

? /:

OpenStudy (zzr0ck3r):

c is what I showed you. notice 1+-isqrt(9) = 1+3i and 1-3i

OpenStudy (anonymous):

Ohhhh! Ok thank you sooo much for all your help!